OpenStudy (anonymous):

A rocket ship travels 14,000 miles per hour. What is the rocket ship's velocity in m/s in correct scientific notation? (1 km = 0.62 mi) A. 6.3 × 10^3 m/s B. 0.627 × 10^3 m/s C. 630 × 10^6 m/s D. 6.3 × 10^5 m/s Please explain?? THanks! :)

OpenStudy (anonymous):

First convert to scientific notation: \(14,000=1.4\times10^4\) \(0.62=6.2\times10^{-1}\) Now some dimensional analysis: \[\frac{1.4\times10^4\text{ mi}}{1\text{ hr}} \times \frac{1\text{ km}}{1\text{ mi}} \times \frac{1\times10^3\text{ m}}{\text{ km}} \times \frac{1\text{ hr}}{6\times10^1\text{ min}} \times \frac{1\text{ min}}{6\times10^1\text{ sec}}\]

OpenStudy (anonymous):

You should notice that the units will cancel to yield \(\dfrac{\text{m}}{\text{sec}}\). Now it's matter of computing the magnitude.

OpenStudy (anonymous):

ohh okay :) so from here, how do you compute the magnitude? :/

OpenStudy (anonymous):

Small typo:\[\frac{1.4\times10^4\text{ mi}}{1\text{ hr}} \times \frac{1\text{ km}}{\color{red}{6.2\times10^{-1}}\text{ mi}} \times \frac{1\times10^3\text{ m}}{\text{ km}} \times \frac{1\text{ hr}}{6\times10^1\text{ min}} \times \frac{1\text{ min}}{6\times10^1\text{ sec}}\]

OpenStudy (anonymous):

ah okay:)

OpenStudy (anonymous):

\[\begin{align*}\frac{1.4\times10^4\times10^3}{6.2\times10^{-1}\times6\times10^1\times6\times10^1}&=\frac{1.4\times10^7}{223.2\times10^1}\\\\ &=\frac{1.4\times10^7}{(2.232\times10^{-2})\times10^1}\\\\ &=\frac{1.4\times10^7}{2.232\times10^{-1}}\\\\ &=\cdots\times10^{6} \end{align*}\] where \(\cdots=\dfrac{1.4}{2.232}\).

OpenStudy (anonymous):

ohh isee :) so 0.627 x 10^6 ? :/

OpenStudy (anonymous):

Right, or \(6.27\times10^5\). If you're accounting for significant digits, you should round up to two: \(6.3\times10^5\).

OpenStudy (anonymous):

ahh okay :) thanks so much!! so basically for problems like this, you keep converting until you hit the scientific notation through simplification?

OpenStudy (e.mccormick):

hmmm....

OpenStudy (anonymous):

Right. It helps to write out each conversion factor like I did in the first comment. You don't have to convert to sci notation right away, it's just a way of easily seeing the powers of 10 reduce.

OpenStudy (e.mccormick):

Did you double check your math Sith?

OpenStudy (anonymous):

ahh okay:) @e.mccormick ?

OpenStudy (anonymous):

I did not :)

OpenStudy (anonymous):

Just double-checked with wolfram, there's a mistake up there. Should be \(6.3\times10^3\) I believe...

OpenStudy (anonymous):

ohhh okay was it just the calculation at the end? :/

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @SithsAndGiggles \[\begin{align*}\frac{1.4\times10^4\times10^3}{6.2\times10^{-1}\times6\times10^1\times6\times10^1}&=\frac{1.4\times10^7}{223.2\times10^1}\\\\ &=\frac{1.4\times10^7}{\color{red}{(2.232\times10^{2})}\times10^1}\\\\ &=\frac{1.4\times10^7}{2.232\times10^{3}}\\\\ &=\cdots\times10^{4} \end{align*}\] \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

ahh okay haha thank you so much!! :D

OpenStudy (anonymous):

Thanks for catching it @e.mccormick

OpenStudy (anonymous):

yes, thank you @e.mccormick !! :D

OpenStudy (e.mccormick):

Way I started it was: \(\dfrac{14000mi}{1h} \cdot \dfrac{1k}{.62mi} = \dfrac{22580.645k}{1h}\) \(\dfrac{22580.645k}{1h} \cdot\dfrac{1h}{60min} = \dfrac{376.344k}{min}\) When you finish with the min to sec and k to meter changes, you get the answer. You can see it is the same sort of thing. As Sith pointed out, one of the keys is watching what cancels: By putting miles on the bottom in the fist conversion, it cancels with the riginal miles. \(\dfrac{14000\bbox[border:2px solid red]{mi}}{1h} \cdot \dfrac{1k}{.62\bbox[border:2px solid red]{mi}} = \dfrac{22580.645k}{1h}\) Same sort of thing with hours, but this time it is on top to cancel what was on the bottom in the original. \(\dfrac{22580.645k}{1\bbox[border:2px solid red]{h}} \cdot\dfrac{1\bbox[border:2px solid red]{h}}{60min} = \dfrac{376.344k}{min}\)

OpenStudy (anonymous):

ahh okay :) so when i do these problems, i'm trying to work on canceling out until i reach the unit/simplification that i'm looking for?

OpenStudy (e.mccormick):

Exactly!

OpenStudy (anonymous):

awesome! I'll be trying to do some more practice problems so i can have this engrained in my brain hehe!! :)

OpenStudy (e.mccormick):

If you are doing it in a calculator the whole way, the method I did works very well. When you are doing it by hand, the method where you trak and cancel \(\times 10^x\) works really well. Not a huge difference between them.

OpenStudy (anonymous):

okay :) thank you!! i'll keep everything in mind (hopefully!!!) :D